Ladt year, Tammy had $30,000 to invest. She invested some of it in an account that paud 7% simple interest per year, and she invested the rest in an account that paid 8% simple interest per year. After 1 year she recueved a total of $2280 in interest. How much did she invest in each account?
Thank you

Answer by richwmiller(17219)

Total amount of money invested: $30000
x+y=30000,
Total yearly interest for the two accounts is: $2280
0.07*x+0.08*y=2280
x=30000-y
Substitute for x
0.07*(30000-y)+0.08*y=2280
Multiply out
2100-0.07*y+0.08*y=2280
Combine like terms.
0.01*y=180
Isolate y
y=$18,000.00 at 8% earns $1,440 interest
x=30000-y
Calculate x
x=$12,000.00 at 7% earns $840 interest
Check
0.07*12000+0.08*18000=2280
840+1440=2280
2280=2280
If 2280=2280 is TRUE and neither x nor y is negative then all is well
codeint1